How do you find the point of intersection between y=-1/2x+4 and y=2x+8
Answer by stanbon(26291)

How do you find the point of intersection between
y=-1/2x+4 and y=2x+8
--------------------------
Substitute for y and solve for "x":
1/2x+4 = 2x+8
---
(2x+4)(2x+8) = 1
4(x+2)(x+4) = 1
4(x^2+6x+8) = 1
x^2+6x+8-1/4
x^2+6x+7.75 = 0
---
x = [-6 +- sqrt(36 -4*7.75)]/2
---
x = [-6 +- sqrt(5))/2
---
x = -3 +- 1.118
---
x = -4.118 or x = -1.882
===========================
Cheers,
Stan H.